a six sided die was rolled 12 times with these results:
Roll- 1, 2, 3, 4, 5, 6
times- 2, 3, 2, 2, 2, 1
What is the experimental probability of rolling an number greater than 2? round as decimal to the nearest hundredth.

Answer by greenestamps(13198)



P = (# of "good" outcomes)/(total # of outcomes)

P(greater than 2) = (2+2+2+1)/(2+3+2+2+2+1) = 7/12